A five character password consists of one letter from the word HELP, followed by four digits which may be 1, 2, 3, 4. If the digits may be used more than once, how many more passwords can be made than if the numbers are used only once?
Answer by sudhanshu_kmr(1152)

no. of ways to select one letter for first place= 4


case 1: if numbers are used only once.
no. of ways to arrange last four digit = 4!

total no. of password possible = 4 * 4! = 96


case 2: if numbers can be repeated

no. of ways to arrange last four digit= 4^4 = 256
total no. of password possible = 4* 256 = 1024


so, (1024 - 96 ) = 928 more passwords possible.....


It is possible that some typing mistake in solution of a problem, if any please ignore it. Understand the concept and try to solve the problem yourself. If there is problem related to concept, contact at
sudhanshu.cochin@yahoo.com or sudhanshu.cochin@gmail.com
Best of luck.......